# Efficient Self-Supervised Learning for Earth Observation via Dynamic Dataset Curation

### v1

@@ Abstract
We release the weights of Nereus-SAR-1, the first foundation model in our Nereus models family — a series of models dedicated to ocean observation and analysis using SAR imagery, at github.com/galeio-research/nereus-sar-models/.
@@ 3.6.2 Comparison to other SAR foundation models
Nereus-SAR-1 ResNet50 75.5 0.75 1.62
Nereus-SAR-1 ViT-S/16 78.6 0.69 1.39
Nereus-SAR-1 ViT-S/8 82.1 0.64 1.38
Nereus-SAR-1 ViT-B/8 83.6 0.63 1.37
@@
Nereus-SAR-1 ResNet50 80.9 0.63 1.33
Nereus-SAR-1 ViT-S/16 89.0 0.57 1.34
Nereus-SAR-1 ViT-S/8 85.8 0.55 1.30
Nereus-SAR-1 ViT-B/8 88.3 0.54 1.29
@@ 4 Discussion and Conclusion
We also address a gap in existing EO foundation models by focusing on a modality—Sentinel-1 WV—for which no pretrained model is well-suited. To support further research, we release a benchmark framework tailored to this modality, relevant for applications ranging from wind field estimation to oil slick detection. We also release the weights for Nereus-SAR-1, the first in a family of models designed specifically for ocean observation using SAR imagery.
@@ 6 Data and Code
The Nereus-SAR-1 model weights are available on HuggingFace at https://huggingface.co/galeio-research/nereus-sar-1. We also release the probing weights for the TenGeoP classification task and the wave height and wind speed regression tasks.

### v2

@@ Abstract
We release the weights of OceanSAR-1, the first foundation model in our OceanSAR models family — a series of models dedicated to ocean observation and analysis using SAR imagery, at github.com/galeio-research/OceanSAR-models/.
@@ 3.6.2 Comparison to other SAR foundation models
OceanSAR-1 ResNet50 75.5 0.75 1.62
OceanSAR-1 ViT-S/16 78.6 0.69 1.39
OceanSAR-1 ViT-S/8 82.1 0.64 1.38
OceanSAR-1 ViT-B/8 83.6 0.63 1.37
@@
OceanSAR-1 ResNet50 80.9 0.63 1.33
OceanSAR-1 ViT-S/16 89.0 0.57 1.34
OceanSAR-1 ViT-S/8 85.8 0.55 1.30
OceanSAR-1 ViT-B/8 88.3 0.54 1.29
@@ 4 Discussion and Conclusion
We also address a gap in existing EO foundation models by focusing on a modality—Sentinel-1 WV—for which no pretrained model is well-suited. To support further research, we release a benchmark framework tailored to this modality, relevant for applications ranging from wind field estimation to oil slick detection. We also release the weights for OceanSAR-1, the first in a family of models designed specifically for ocean observation using SAR imagery.
@@ 6 Data and Code
The OceanSAR-1 model weights are available on HuggingFace at https://huggingface.co/galeio-research/OceanSAR-1. We also release the probing weights for the TenGeoP classification task and the wave height and wind speed regression tasks.